OpenStudy (anonymous):

What is the measure of angle SQR? Given: Angle SQR: (x-1) Angle SQP: (2x+4) Its a 90 degree angle. I just need to find the measure of angle SQR. The answer options are : 45 , 62, 28 , or 29 .

OpenStudy (anonymous):

we have : x-1+2x+4=90 Right ? @loe1122

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

So we have : x-1+2x+4=90 2x=90-4+1 2x=87 x=87/2 Then : SQR : x-1 => 87/2 + 1 =>43.5-1=42.5 Got it ?:)

OpenStudy (anonymous):

But thats not one of the answers lol

OpenStudy (anonymous):

Is that equation I said is for your question ?